# Co-Paced Learning Strategy Based on Confidence for Flying Bird Object Detection Model Training

### v1

@@ 5 Experiments
To verify that the CPL-BC strategy proposed in this paper is effective and advanced in the training process of the FBOD model for surveillance video, we designed and carried out quantitative and qualitative comparative experiments. Specifically, we compare four other model training strategies: All Sample training strategy (AS, All Samples are used, and the model is trained by general training strategy that do not distinguish the difficulty of samples), Easy Sample training strategy (ES, Select Easy samples and train the model with general training strategy), loss-based Self-Paced Learning strategy [17, 18], Self-Paced Learning strategy Based on Confidence (SPL-BC) [12]. In the loss-based Self-Paced Learning strategy, three regularizers, Hard [17], Linear [18], and Logarithmic [18], are used in comparison, which are respectively recorded as SPL-BH, SPL-BLine, and SPL-BLog. For specific implementation details and parameter settings of all comparison training strategies, please refer to the relevant part of the literature [12].
@@ 5.4.1 The Auantitative Experiment
The quantitative experimental results of the detection performance of the FBOD models trained with various training strategies on the FBD-SV-TSS dataset and tested on the test set are shown in Table 2. The experimental results indicate that the FBOD model trained using the CPL-BC strategy (ESP) proposed in this paper achieves an $\text{AP}_{50}$ of 79.0% on the FBD-SV-TSS dataset’s test set. Specifically, this strategy improves the $\text{AP}_{50}$ by 2.8% compared to the AS strategy, 6.8% compared to the ES strategy, and 0.7% compared to the SPL-BLine [18] training strategy, which performs relatively best among the loss-based SPL strategies. Additionally, it improves by 0.8% compared to the SPL-BC [12] strategy. Furthermore, this strategy’s False Detection Rate (FDR) is reduced by 6.5% compared to the AS strategy but is slightly higher than that of the SPL-BC [12] strategy. It is noteworthy that training the FBOD model with the CPL-BC strategy using a Prior based on Easy Samples (ESP) results in an $\text{AP}_{50}$ that is only 0.2% higher than that of using a Prior based on the All Samples (ASP).

### v2

@@ 5 Experiments
To verify that the CPL-BC strategy proposed in this paper is effective and advanced in the training process of the FBOD model for surveillance video, we designed and carried out quantitative and qualitative comparative experiments. Specifically, we compare four other model training strategies: All Sample training strategy (AS, All Samples are used, and the model is trained by general training strategy that do not distinguish the difficulty of samples), Easy Sample training strategy (ES, Select Easy samples and train the model with general training strategy), loss-based Self-Paced Learning strategy [15, 16], Self-Paced Learning strategy Based on Confidence (SPL-BC) [10]. In the loss-based Self-Paced Learning strategy, four regularizers, Hard [15], Linear [16], Logarithmic [16], and Polynomial [23] are used in comparison, which are respectively recorded as SPL-BH, SPL-BLine, SPL-BLog, and SPL-BPoly. For specific implementation details and parameter settings of all comparison training strategies, please refer to the relevant part of the literature [10].
@@ 5.4.1 The Auantitative Experiment
The quantitative experimental results of the detection performance of the FBOD models trained with various training strategies on the FBD-SV-TSS dataset and tested on the test set are shown in Table 2. The experimental results indicate that the FBOD model trained using the CPL-BC strategy (ESP) proposed in this paper achieves an $\text{AP}_{50}$ of 79.0% on the FBD-SV-TSS dataset’s test set. Specifically, this strategy improves the $\text{AP}_{50}$ by 2.8% compared to the AS strategy, by 6.8% compared to the ES strategy, and by 0.7% compared to the SPL-BLine [16] training strategy, which performs relatively best among the loss-based SPL strategies. Additionally, it improves by 0.8% compared to the SPL-BC [10] strategy. Furthermore, this strategy’s False Detection Rate (FDR) is reduced by 6.5% compared to the AS strategy but is slightly higher than that of the SPL-BC [10] strategy. It is noteworthy that training the FBOD model with the CPL-BC strategy using a Prior based on Easy Samples (ESP) results in an $\text{AP}_{50}$ that is only 0.2% higher than that of using a Prior based on the All Samples (ASP).